OpenStudy (anonymous):

plz help me with an 8th grade question An equation is shown below: 6(3x − 7) = 2 Which of the following correctly shows the beginning steps to solve this equation? Step 1: 9x − 1 = 2 Step 2: 9x = 3 Step 1: 9x − 7 = 2 Step 2: 9x = 9 Step 1: 18x − 7 = 2 Step 2: 18x = 9 Step 1: 18x − 42 = 2 Step 2: 18x = 44

OpenStudy (anonymous):

@Agl202 @ikram002p @Harley1312 @nazeeh @Xnlovely @moneymob2434 @kidrah69 @redbeardd

OpenStudy (harley1312):

@selfconfidence777

OpenStudy (anonymous):

@emilyfuchs @dmndlife24 @DarkMoonZ @deana99 @DARTHVADER2900 @elizadaniel @Ellie202000 @Ellie202000 @Ellen_mariexo @QueenEmilyofHearts @welshfella @willsgirl16 @willman8 @Australopithecus @Agl202 @AJ01 @angel12310 @Firefly007 @FallenAngelQueen @FallenAngelQueen @Forever_Malia @gigas @GreatGuy @GoArmy @gibsonboy11

OpenStudy (anonymous):

plz help

OpenStudy (darthvader2900):

Please do not mass tag.

OpenStudy (selfconfidence777):

step 1

OpenStudy (anonymous):

D

OpenStudy (anonymous):

ok sorry i just really need help this is my last question and i only have 5 minutes to finish

OpenStudy (selfconfidence777):

18x-7=2

OpenStudy (anonymous):

6(3x-7)=2 6*3x-6*7=2 18x-42=2

OpenStudy (anonymous):

so its d

OpenStudy (anonymous):

4th 18x-42=2 18x=42+2 18x=44

OpenStudy (anonymous):

yep rigght

OpenStudy (anonymous):

yes

OpenStudy (selfconfidence777):

i think its C lol

OpenStudy (anonymous):

ok thanks